solve using elimination method
x+3y=-20
4x+3y=-17

Answer by mananth(16946)

x+3y=-20 .............1
4x+3y=-17 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-x-3y=20
4x+3y=-17
Add the two equations
3x = 3
/3
x=1
plug value of x in (1)
x+3y=-20
1+3y=-20
3y=-20-1
3y=-21
y=-7
unique solution ( 1,-7)